A 10-kg metal ball is dropped from the top of a building 5.0 - m high. How long will it take this object to hit the ground? Neglect air resistance.


1
Expert's answer
2021-11-15T10:33:15-0500

"v_0 = 0"

"s = 5 m"

"g= 9.81\frac{m}{s^2}"

"s = v_0t+\frac{gt^2}{2}"

"t = \sqrt{\frac{2s}{g}}= \sqrt{\frac{10}{9.81}}\approx1s"


"\text{Answer: }1s"